It is now December 2017. The current interest rate is 2%. The May 2018 futures price for gold is $1,500, whereas the December 2018 futures price is $1510. Is here an arbitrage possibility here? If so, how would you exploit it?
Answer #1

Answer:

The spot price implied by the December futures price is

= 1510 / [1 + 0.02] = 1,480.39

which is greater than the spot price implied by the May futures price:

1500 / [1 + 0.02]^½ = 1500 / 1.01 = 1,485.15

So since the actual December futures price is too low, the arbitrage involves selling the May futures contract and buying the December futures contract.